# Neutral Barium in Solid Neon: Optical Spectroscopy and First Excited State Lifetime

### v1

@@ III.3 Lifetime of first excited state 5d6s 3D1
As shown in the previous sections, and schematically in Figure 4, a single laser allows us to populate the 5d6s 3Dj manifold, similar to the earlier work of Lebedev et al. Weis_BaHelium, where the 5d6s 3D1 state of barium in a helium matrix at 1.5 K is populated using the pulsed second harmonic of a Nd:YAG laser at approximately 18 80018\,800 cm-1 .
In our experiments with barium in a solid neon at 6.8 K, this state is populated using continuous wave laser light tuned to 13 514 cm-1, enabling a direct comparison with the lifetime measurement performed in solid helium and in the vacuum gas phase.
@@
We obtain a lifetime more than two orders of magnitude higher than the gas phase measurement, which is estimated to be 1.5 $\pm$ 0.2 ms AdamsLeiHusain+2001+1205_lifetimeBaGas. This increase might indicate the weak interaction of the host matrix, as neon low phonon density and weak interaction with the embedded atoms enable long lived excited states. For a comparison, in solid helium, an even better lifetime of $2.72\pm 0.04$ s is observed at a temperature of 1.5 K Weis_BaHelium.
@@
A rate equation model has been used to fit the obtained lifetime and extrapolate the possible lifetime values at lower temperatures. It considers both radiative ($\tau_{r}$) and non-radiative ($\tau_{nr}$) decay channels, described by the two termsfit_description_PL, torVergata_fitArrhenius:
@@
Fit results are shown in Table 2, where $E_{a}$ is the activation energy . The lifetime is expected to increase to approximately 0.42 $\pm$ 0.03 s at 2 K.
@@ IV Conclusions
For the time in literature, we observed barium atoms embedded in a neon matrix through a combination of cascade florescence and selective excitation using a third harmonic 1064 nm pulsed laser and single or double continuous wave lasers, exploiting long living metastable state of barium within the matrix.
@@
Our lifetime measurement serves as a reference point for future studies aiming to explore similar systems in parahydrogen matrices. Previous work by Lancaster et al. Lancaster_2021_NevsPh found no detectable fluorescence for rubidium in parahydrogen, whereas clear emission was observed for boron Tam2000_boron_pH and other molecules Huang2016_molecules_pH. A similar investigation for Ba and BaF will be crucial to establish parahydrogen as a next-generation host for matrix-isolated spectroscopy and precision measurements of fundamental properties.

### v2

@@ III.3 Lifetime of first excited state 5d6s 3D1
As shown in the previous sections, and schematically in Figure 4, a single laser allows us to populate the 5d6s 3Dj manifold, similar to the earlier work of Lebedev et al. Lebedev et al. [2011], where the 5d6s 3D1 state of barium in a helium matrix at 1.5 K is populated using the pulsed second harmonic of a Nd:YAG laser at approximately 18 80018\,800 cm-1 .
In our experiments with barium in a solid neon at 6.8 K, this state is populated using continuous wave laser light tuned to 13 514 cm-1, enabling a direct comparison with the lifetime measurement performed in solid helium.
@@
For a comparison, in solid helium, a longer lifetime of $2.72\pm 0.04$ s is observed at a temperature of 1.5 K Lebedev et al. [2011].
@@
A rate equation model has been used to fit the obtained lifetime and extrapolate the possible lifetime values at lower temperatures. It considers both radiative ($\tau_{r}$) and non-radiative ($\tau_{nr}$) decay channels, described by the two termsKobitski et al. [2003], Prestopino et al. [2019]:
@@
The fit results are shown in Table 2, where $E_{a}$ is the activation energy. The lifetime is expected to increase to approximately 0.42 $\pm$ 0.03 s at 2 K. T The model captures the expected low temperature plateau corresponding to a regime where the dopant non radiative decay channels are suppressed. Moreover, in the precedent work of BaF in neonLi et al. [2023], a similar model has been used, with a resulting activation energy value of $E_{a(BaF:Ne)}=hc\times 16.2(3)\;\text{cm${}^{-1}$}=(2.01\pm 0.04)\times 10^{-3}\;\text{eV}$, consistent with the one obtained in this work, shown in shown in Table 2. Future work and new models might take into account additional mechanisms inherent to matrix environments, such as site dependent trapping, crystal inhomogeneity and local stress of laser excitation light.
@@ IV Conclusions
For the first time in literature, we observed barium atoms embedded in a neon matrix through a combination of cascade fluorescence and selective excitation using a third harmonic 1064 nm pulsed laser and single or double continuous wave lasers, exploiting long living metastable state of barium within the matrix.
@@
Our lifetime measurement serves as a reference point for future studies aiming to explore similar systems in parahydrogen matrices. Previous work by Lancaster et al. Lancaster et al. [2021] found no detectable fluorescence for rubidium in parahydrogen, whereas clear emission was observed for boron Tam et al. [2000] and other molecules Huang and Lee [2016]. A similar investigation for Ba and BaF will be crucial to establish parahydrogen as a next-generation host for matrix-isolated spectroscopy and precision measurements of fundamental properties.